# State-space kinetic Ising model reveals task-dependent entropy flow in sparsely active nonequilibrium neuronal dynamics

### v2

@@ 3.2 Independent neurons
Here we consider independent neurons (i.e., no couplings $theta_{ij}=0$) with time-varying field $\theta_{i,t}$. The entropy flow in this system is caused solely by the time-varying fields, or equivalently, the activity rate of individual neurons.
@@
$g_{i,t,s}$ $=\theta_{i,t},$ (176)
$\Delta_{i,t,s}$ $=0,$ (177)
which is indpendent of $s$, which once again makes the inputs to $\chi$ and $\phi_{i}$ common. Then, we have
@@
$=\sum_{i}\left(m_{i,t}-m_{i,t-1}\right)\cdot\theta_{i,t}$ (178)
For $\theta_{i,t}<0$, which corresponds to $m_{i,t}<0.5$, decrease in the activity rate $m_{i,t}-m_{i,t-1}<0$ yields positive entropy flow, and increase in the activity rate induces negative entropy flow.
4 S4 text. The d-prime measure
@@
To evaluate the sensitivity of the mice to the stimulus, the primary behavioral metric, $d^{\prime}$ (d-prime), was calculated using data detected only in the active condition with visual changes. The formula for $d^{\prime}$ is as follows:
$d^{\prime}=Z(R_{H})-Z(R_{F}),$ (179)
@@
$\frac{1}{2N_{H}}\leq R_{H}\leq 1-\frac{1}{2N_{H}},\quad\frac{1}{2N_{F}}\leq R_{F}\leq 1-\frac{1}{2N_{F}},$ (180)
@@
5 S5 text. Entropy flow of high-firing neurons and behavioral performance
@@
To elucidate how individual neurons increase total entropy flow in the active condition despite that a smaller fraction of neurons exhibits substantial firing rates (S1 Fig. S2 Fig., and S3 Fig.), we examined the relationship between the entropy flow and spike rates of individual neurons.
@@
As shown in Eqs. 13 and 14, the mean-field entropy flow can be decomposed into contributions from individual neurons. We computed the entropy flow of individual neurons under the active and passive conditions and compared them with their firing rates (Fig. S8 FigA, mouse 574078). The dotted lines connect the values for the active (red) and passive (blue) conditions. We then investigated whether the change in the entropy flow by the behavioral conditions depends on the neuron’s firing rate. Figure S8 FigB shows the relationship between the geometric mean spike rates of the two conditions (abscissa) and the difference in entropy flow (ordinate) for individual neurons. The difference was computed as ‘active’ - ‘passive,’ indicating that the positive value marks a larger entropy flow in the active condition. The positive Spearman rank correlation coefficient ($\rho=0.22$) for this exemplary mouse suggests the tendency that neurons with higher spike rates contributed to increasing total entropy flow in the active condition, despite the summed entropy flow differences across all individual neurons being negative (3.8331-3.8331 for this mouse). However, significant variations in the rank correlations were observed across mice.
@@
Assuming that fewer high-firing neurons in the sparsely active populations in the active condition play a critical role in sensory processing (i.e., sparse coding [83, 84, 85]) and that such sensory processing involves time-asymmetric causal patterns, we hypothesized that the above relationship between the spike rates and entropy flow change might be related to mice’s cognitive performance. To evaluate the task sensitivity of the mice, we used the primary behavioral metric, $d^{\prime}$ (mean d-prime, see Supporting Information S4 text for its definition). The scatter plot in Fig. S9 FigA illustrates the relationship between behavioral measures (mean d-prime) and the rank correlation of entropy flow change with spike rates for all mice for image ’im036_r’. The plot suggests a positive dependency between these two values ($\rho=0.3578$ measured by the Spearman rank correlation). To confirm this result, we conducted the permutation test that compared the observed rank correlation of the scatter plot with those of the surrogate data constructed by permuting the values of mean d-prime (Fig. S9 FigA Right). The result confirms the statistical significance of the positive correlation ($p=0.0304$).
@@
To corroborate that the result does not reflect estimation error in couplings, we analyzed trial-shuffled data, which showed no clear trend (Fig. S9 FigB). A permutation test confirmed that the observed correlation yielded a non-significant p-value of 0.5063. This result confirms that the association between higher entropy flow and higher firing neurons in more task-sensitive mice was driven by significant changes in the coupling strengths between the active and passive conditions, rather than firing rate shifts or noise couplings.
@@
S1 Fig: Spike-rate dynamics and distributions for mice 1-18 Spike-rate dynamics and distributions under the active (red) and passive (blue) conditions. The presentation styles for each mouse follow Fig. 6A. The mice were listed in descending order of behavioral performance measured by d-prime. See Fig. S2 Fig for the remaining mice.
S2 Fig: Spike-rate dynamics and distributions for mice 19-37 The same as in S1 Fig but for the remaining 19 mice.
S3 Fig: Comparison of mean spike rates and coefficient of variation in the active and passive conditions. A Mean spiking probability across all bins, trials, and neurons in active and passive conditions. Each line represents the same mouse. Neurons showed significantly lower firing rates in the active condition ($p=1.556\times 10^{-8}$, Wilcoxon signed-rank test). B Coefficient of variations (CVs) of the firing rate distributions, a measure of sparseness, in the active and passive conditions. CV was significantly higher in the active condition ($p=8.35\times 10^{-8}$, Wilcoxon signed-rank test).
S4 Fig: Estimated neural dynamics under active and passive conditions in shuffled data of mouse 574078. Presentation style follows Fig. 6.
S5 Fig: Mean effective coupling and shuffle control. A Violin plots of time-averaged effective couplings for mouse $574078$ under the active and passive conditions. Horizontal bars indicate the mean (red) and median (green); points show individual matrix entries. B Population summary of the per-mouse mean coupling in the original data; each line connects the active and passive values from the same mouse. Panel annotations report $p$-values and sample size ($n$) from Wilcoxon signed-rank tests across mice (two-sided). C Shuffle-adjusted means, where for each mouse the value in each condition is computed as (Original - Shuffle); lines connect paired values. The annotation reports a one-sided Wilcoxon signed-rank test assessing whether the median of {(Original - Shuffle) in Active} minus {(Original - Shuffle) in Passive} is greater than zero. Together, the results indicate that the mean effective coupling is larger in the active condition than in the passive condition, and that this increase persists after shuffle correction.
S6 Fig: Time courses of entropy flow and mean spike rates for each mouse under active and passive conditions. Each subplot represents the dynamics of an individual mouse. Solid lines are entropy flows (red for active, blue for passive) while dashed lines represent the average population spike rate (red for active, blue for passive).
@@
S8 Fig: Relationship between spike rates and entropy flow for an example mouse. A Mean spike rate vs entropy flow per individual unit under the active and passive conditions (mouse 574078). Dashed lines connect values for the two conditions, highlighting behavioral state-dependent changes. B Geometric mean spike rate (abscissa) vs differences in entropy flow (ordinate) for individual units. The positive Spearman correlation coefficient ($\rho=0.22$) suggests that units with higher spike rates increased entropy flow in the active condition.
S9 Fig: Relating the dependency of entropy flow change of individual units on firing rates with behavioral performance. A (Left) Scatter plot of behavioral performance (mean d-prime) vs. the Spearman rank correlation between the geometric mean rate and entropy flow change of individual units (see Fig. S8 Fig). Each dot represents a single mouse. The dependency in this scatter plot was assessed again by the Spearman rank correlation coefficient, yielding $\rho=0.3578$. (Right) A permutation test comparing the observed correlation value $\rho$ with those obtained from the surrogate data. A statistically significant positive relationship was observed ($p=0.0304$, two-tailed). The surrogate data was constructed by permuting the values of mean d-prime. B Results for trial-shuffled data.

### v3

@@ 3.2 Independent neurons
Here we consider independent neurons (i.e., no couplings $\theta_{ij}=0$) with time-varying field $\theta_{i,t}$. The entropy flow in this system is caused solely by the time-varying fields, or equivalently, the activity rate of individual neurons.
@@
$g_{i,t,s}$ $=\theta_{i,t},$ (S3.94)
$\Delta_{i,t,s}$ $=0,$ (S3.95)
which is independent of $s$, making the inputs to $\chi$ and $\phi_{i,t}$ common once again. Then, we have
@@
$=\sum_{i}\left(m_{i,t}-m_{i,t-1}\right)\cdot\theta_{i,t}.$ (S3.96)
For $\theta_{i,t}<0$, which corresponds to $m_{i,t}<0.5$, a decrease in the activity rate $m_{i,t}-m_{i,t-1}<0$ yields positive entropy flow, and an increase in the activity rate induces negative entropy flow.
Supplementary Note 4 The d-prime measure
@@
To evaluate the sensitivity of the mice to the stimulus, the primary behavioral metric, $d^{\prime}$, was calculated using data detected only in the active condition with visual changes. The formula for $d^{\prime}$ is as follows:
$d^{\prime}=Z(R_{H})-Z(R_{F}),$ (S4.97)
@@
$\frac{1}{2N_{H}}\leq R_{H}\leq 1-\frac{1}{2N_{H}},\quad\frac{1}{2N_{F}}\leq R_{F}\leq 1-\frac{1}{2N_{F}},$ (S4.98)
@@
Supplementary Note 5 Entropy flow of high-firing neurons and behavioral performance
@@
To elucidate how individual neurons increase total entropy flow in the active condition despite a smaller fraction of neurons exhibiting substantial firing rates (Supplementary Fig. S2, Fig. S3, and Fig. S4), we examined the relationship between the entropy flow and spike rates of individual neurons.
@@
As shown in Eqs. 13 and 14, the mean-field entropy flow can be decomposed into contributions from individual neurons. We computed the entropy flow of individual neurons under the active and passive conditions and compared them with their firing rates (Supplementary Fig. S8A, mouse 574078). The dotted lines connect the values for the active (red) and passive (blue) conditions. We then investigated whether the change in the entropy flow by the behavioral conditions depends on the neuron’s firing rate. Supplementary Fig. S8B shows the relationship between the geometric mean spike rates of the two conditions (abscissa) and the difference in entropy flow (ordinate) for individual neurons. The difference was computed as ‘active’ - ‘passive,’ indicating that the positive value marks a larger entropy flow in the active condition. The positive Spearman rank correlation coefficient ($\rho=0.22$) for this exemplary mouse suggests that neurons with higher spike rates contributed to increasing total entropy flow in the active condition, despite the summed entropy flow differences across all individual neurons being negative (3.8331-3.8331 for this mouse). However, significant variations in the rank correlations were observed across mice.
@@
Assuming that fewer high-firing neurons in the sparsely active populations in the active condition play a critical role in sensory processing (i.e., sparse coding [4, 5, 6]) and that such sensory processing involves time-asymmetric causal patterns, we hypothesized that the above relationship between the spike rates and entropy flow change might be related to mice’s cognitive performance. To evaluate the task sensitivity of the mice, we used the primary behavioral metric, $d^{\prime}$ (mean d-prime, see Supplementary Note 4 for its definition). The scatter plot in the left panel of Supplementary Fig. S8C illustrates the relationship between behavioral measures (mean d-prime) and the rank correlation of entropy flow change with spike rates for all mice for image ’im036_r’. The plot suggests a positive dependency between these two values ($\rho=0.3578$ measured by the Spearman rank correlation). To confirm this result, we conducted the permutation test that compared the observed rank correlation of the scatter plot with those of the surrogate data constructed by permuting the values of mean d-prime (Supplementary Fig. S8C Right). The result confirms the statistical significance of the positive correlation ($p=0.0304$).
@@
To corroborate that the result does not reflect estimation error in couplings, we analyzed trial-shuffled data, which showed no clear trend (Supplementary Fig. S8D). A permutation test confirmed that the observed correlation yielded a non-significant p-value of 0.5063. This result confirms that the association between higher entropy flow and higher firing neurons in more task-sensitive mice was driven by significant changes in the coupling strengths between the active and passive conditions, rather than firing rate shifts or noise couplings.
@@
Supplementary Fig. S3: Spike-rate dynamics and distributions for mice 19-37. The same as in Supplementary Fig. S2 but for the remaining 19 mice.
Supplementary Fig. S4: Comparison of mean spiking probability and coefficient of variation in the active and passive conditions. A Mean spiking probability across all bins, trials, and neurons in active and passive conditions. Each line represents the same mouse. Neurons showed significantly lower firing rates in the active condition ($p=1.556\times 10^{-8}$, Wilcoxon signed-rank test). B Coefficient of variations (CVs) of the firing rate distributions, a measure of sparseness, in the active and passive conditions. CV was significantly higher in the active condition ($p=8.35\times 10^{-8}$, Wilcoxon signed-rank test).
Supplementary Fig. S5: Mean effective coupling and shuffle control. A Violin plots of time-averaged effective couplings for mouse $574078$ under the active and passive conditions. Horizontal bars indicate the mean (red) and median (green); points show individual entries. B Population summary of the per-mouse mean coupling in the original data; each line connects the active and passive values from the same mouse. Panel annotations report $p$-values and sample size ($n$) from Wilcoxon signed-rank tests across mice (two-sided). C Shuffle-adjusted means, where for each mouse the value in each condition is computed as (Original - Shuffle); lines connect paired values. The annotation reports a one-sided Wilcoxon signed-rank test assessing whether the median of {(Original - Shuffle) in Active} minus {(Original - Shuffle) in Passive} is greater than zero. Together, the results indicate that the mean effective coupling is larger in the active condition than in the passive condition, and that this increase persists after shuffle correction.
Supplementary Fig. S6: Time courses of entropy flow and mean spike rates for each mouse under active and passive conditions. Each subplot represents the dynamics of an individual mouse. Solid lines are entropy flows (red for active, blue for passive) while dashed lines represent the average population spike rate (red for active, blue for passive).
@@
Supplementary Fig. S8: Relating the dependency of entropy flow change of individual units on firing rates with behavioral performance. A Mean spike rate vs entropy flow per individual unit under the active and passive conditions (mouse 574078). Dashed lines connect values for the two conditions, highlighting behavioral state-dependent changes. B Geometric mean spike rate (abscissa) vs differences in entropy flow (active - passive, ordinate) for individual units. The positive Spearman correlation coefficient ($\rho=0.22$) suggests that units with higher spike rates increased entropy flow in the active condition. C (Left) Scatter plot of behavioral performance (mean d-prime) vs. the Spearman rank correlation between the geometric mean rate and entropy flow change of individual units. Each dot represents a single mouse. The dependency in this scatter plot was assessed again by the Spearman rank correlation coefficient, yielding $\rho=0.3578$. (Right) A permutation test comparing the observed correlation value $\rho$ with those obtained from the surrogate data. A statistically significant positive relationship was observed ($p=0.0304$, two-tailed). The surrogate data was constructed by permuting the values of mean d-prime. D Results for trial-shuffled data.